the car increased its speed from 86.4 km / h to 133.2 km / h in 2.5 seconds.
Calculations of the acceleration of the car!
1
Expert's answer
2021-02-28T07:38:50-0500

Initial speed of the car(u) = 86.4 km/h = 24 m/s

Final speed of the car(v) = 133.2 km/h = 37 m/s

Time taken = 2.5 s

"a=\dfrac{v-u}{t}=\dfrac{37-24}{2.5}"

"\Rightarrow a=5.2m/s^2"